OpenStudy (anonymous):

1000(0.98) to the power of N = 250 (0.98)to the power of N = 1/4 0.98)= 1/4 to the power of 1/N please help me solve WITHOUT using LOG

OpenStudy (anonymous):

using Log the answer is 68.6 how do i get this with out using log?

OpenStudy (phi):

I know there are approximations to the ln, if that is what you mean: \[ ln(1±x) \approx ±x -\frac{x^2}{2} ±\frac{x^3}{3} \] you can use that, and (knowing?) ln(2)= 0.693 (some people memorize this!) N ln(1-0.02)= ln(2^-2) N ( -0.02 - 0.0004/2 ) = -2*0.693 (ignore x^3 term) N= -2*0.693/-0.0202 = 68.6

OpenStudy (anonymous):

thankyou!